A pair of 6-sided dice are rolled. What is the probability that the sum of the two numbers on the dice will be 5?
1
Expert's answer
2011-10-13T10:12:46-0400
The total number of possible cases is 6*6 = 36. Among them, the are 4 successful ones: when the dice rolls 1 and 3, 2 and 3, 3 and 2, 4 and 1. So, the probability is 4/36 = 1/9.